OpenStudy (anonymous):

Terry types twice as many words per minute as Cecilia. Paige types 5 words per minute less than Terry. Reggie types 10 words per minute more than Terry. How many words per minute does each person type if the total number of words per minute that Reggie and Cecilia type equals the total number of words per minute that Terry and Paige type?

OpenStudy (campbell_st):

well let cecilia's speed by x and use this to find the speed of everyone Cecilia = x Terry = 2x Reggie 2x + 10 Paige = 2x - 5 so using the information you can equate things Reggie + Cecilia = Terry and Paige 2x + 10 + x = 2x + 2x - 5 or simply 3x + 10 = 4x - 5 now solve for x. hope it helps